# section[1.5em]1.5em1pc

### v2

@@ 5.   Beck modules over $A$
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right)=\\ \sum_{t+t^{\prime}=nq_{j},t^{\prime}>0}\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),\left((n)\diamond({\underline{q}})\right)\circ_{j}(t,t^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u}).$
@@
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{1},\dots,a_{u},m_{u})\right)=\\ \sum_{t+t^{\prime}=nq_{u},t^{\prime}>0}\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),\left((n)\diamond({\underline{q}})\right)\circ_{j}(t,t^{\prime})}(a_{1},\dots,a_{u},m_{u}).$
@@
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{1},\dots,a_{u},m_{u})\right)=\\ \sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma}x(y^{\times n}),(\lambda,\lambda^{\prime})\diamond({\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime}))}(a_{1},\dots,a_{u},a_{1},\dots,a_{u},m_{u}),$
@@
$(R_{\lambda,\lambda^{\prime}}^{l,l^{\prime}})_{u}=\{k-q_{u}+\alpha k+\gamma,\alpha\in[\lambda],\gamma\in[q_{u}]\}\cup\{\lambda k+(k-q_{u})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l]\}.$
$(R_{\lambda,\lambda^{\prime}}^{l,l^{\prime}})_{u+1}=\{\lambda k+(k-l^{\prime})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l^{\prime}]\}.$
@@
$\sum_{\sigma^{\prime\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\Sigma_{(\lambda,\lambda^{\prime})\diamond({\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime}))}}\sigma\sum_{\Sigma_{(\lambda,\lambda^{\prime})\diamond({\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime}))}/\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma x(y^{\times n})=\\ \sum_{\sigma^{\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma^{\prime}x(y^{\times n}),$
@@
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{1},\dots,a_{u},m_{u})\right)=\\ \sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma^{\prime}}\sigma^{\prime}x(y^{\times n}),R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}(a_{1},\dots,a_{u},m_{u}).$
@@
Assigning to each $A$-module $M$ the Beck module $A\ltimes M$ obtained by equipping the vector space $A\oplus M$ with the divided power $\mathcal{P}$-algebra structure given by (**∗ ∗5) provides a functor:
@@
Abelian group objects in $\Gamma(\mathcal{P})$-algebras are equivalent to abelian $\Gamma(\mathcal{P})$-algebras from Definition 4.1, that is:
@@
Applying 5.2 in the case $A=0$, the data of an abelian group object in $\Gamma(\mathcal{P})$-algebras is equivalent to the data of a module over the $\Gamma(\mathcal{P})$-algebra 00. Such a 00-module $M$ is an abelian $\Gamma(\mathcal{P})$-algebra equipped with operations $\beta_{x,(r_{1},\dots,r_{s})}(0,\dots,0,-)\colon M\to M$. If $s>1$ and if at least one of the $r_{i}$ for $i<s$ is non-zero, then relation ($\beta$A3) implies that $\beta_{x,(r_{1},\dots,r_{s})}(0,\dots,0,m)=0$ for all $M$. If $s>1$ and $r_{i}=0$ for all $i<s$, relation ($\beta$A2) shows that $\beta_{x,(r_{1},\dots,r_{s})}(0,\dots,0,m)=\beta_{x,(r_{s})}(m)$. Finally, when $s=1$, relation ($\beta$AM7) shows that $\beta_{x,(r_{s})}(m)=\beta_{x}(m)$ is provided by the structure of abelian $\Gamma(\mathcal{P})$-algebra on $M$. Thus, a 00-module is an abelian $\Gamma(\mathcal{P})$-algebra equipped with additionnal operations which are all trivial. ∎
@@ 6.   Universal enveloping algebra
    $\beta_{\lambda x+y,{\underline{r}}}(a_{1},\dots,a_{s-1},-)-\left(\lambda\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)+\beta_{y,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\right)$ for all $x,y\in\mathcal{P}(n)^{{\underline{r}}}$, $\lambda\in\mathbb{F}$,
@@
    where ${\underline{q}}^{\prime}=({\underline{q}},0)$,
@@
    $\beta_{x,{\underline{r}}}(a_{0}+a_{1},\dots,a_{s-1},-)-\left(\sum_{l+l^{\prime}=r_{1}}\beta_{x,{\underline{r}}\circ_{1}(l,l^{\prime})}(a_{0},a_{1},\dots,a_{s-1},-)\right)$,
@@
    $\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)$ for all $x$ such that there exists $q_{1}+\dots+q_{u}=r_{s}$, $y\in\mathcal{P}(n)^{\Sigma_{{\underline{r}}\circ_{s}{\underline{q}}}}$, such that $x=\sum_{\sigma\in\Sigma_{n-r_{s}}\times\Sigma_{r_{s}}/\Sigma_{n-r_{s}}\times\Sigma_{{\underline{q}}}}\sigma y$, and at least two of the $q_{j}$s are non-zero.
@@
Following 4.6, we will allow the notation:
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},-,a_{i+1},\dots,a_{s}):=\beta_{\sigma_{i}^{*},{\underline{r}}^{\sigma_{i}}}(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{s},-).$
@@
$\mu:\mathbb{U}_{\Gamma(\mathcal{P})}(A){}_{\mathbb{F}}\!\otimes_{\mathbb{F}}\mathbb{U}_{\Gamma(\mathcal{P})}(A)\to\mathbb{U}_{\Gamma(\mathcal{P})}(A)$
@@
The multiplication $\mu$ endows $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ with the structure of an associative unital ring.
@@
Let $V$ be the $1$-dimensional $\mathbb{F}$-vector space spanned by $v$. Consider the $\Gamma(\mathcal{P})$-algebra $B$ obtained as a quotient of the $\Gamma(\mathcal{P})$-algebra $\Gamma(\mathcal{P},A\oplus V)$ by the $\Gamma(\mathcal{P})$-ideal generated by the two following family of elements:
@@
We will see a concrete example with the enveloping algebra $V(A)=\mathbb{U}_{\Gamma(\mathrm{Com})}(A)$ in Section 9. In the special case $\mathbb{F}=\mathbb{F}_{p}$, the equality $\lambda^{r_{s}}=\lambda$ holds (since $r_{s}$ is a power of $p$), and thus $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ is an $\mathbb{F}$-algebra.
@@
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{j-1},-,a_{j+1},\dots,a_{s}):=\beta_{\sigma_{j}x,{\underline{r}}^{\sigma_{j}^{*}}}(a_{1},\dots,a_{j-1},a_{j+1},\dots,a_{s},-)\in\mathbb{U}_{\Gamma(\mathcal{P})}(A).$
@@
Denote by $\mathbb{U}_{\mathrm{Ab}}(\Gamma(\mathcal{P}))\subset\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ the vector subspace spanned by the elements $\beta_{x,(n)}(-)$. Note that $\mathbb{U}_{\mathrm{Ab}}(\Gamma(\mathcal{P}))$ is a unital associative subring and does not depend on $A$.
@@
$\beta_{x}(m):=\beta_{x,(n)}(-)\cdot m.$
@@
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},m):=\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\cdot m.$
@@
Let $M$ be an abelian $\Gamma(\mathcal{P})$-algebra. Then, $\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes_{\mathbb{U}_{\mathrm{Ab}(\Gamma(\mathcal{P}))}}M$ is the free $A$-module generated by $M$.
@@
Note that here, in the expression $\mathbb{U}_{\mathcal{P}}(A)$, we have dropped the forgetful functor $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\mathcal{P}\text{-}\mathrm{Alg}$ which is induced by the norm map as well.

### v3

@@ 5.   Beck modules over $A$
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\bigg{(}\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\bigg{)}=\\ \sum_{t+t^{\prime}=nq_{j},t^{\prime}>0}\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),\left((n)\diamond({\underline{q}})\right)\circ_{j}(t,t^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u}).$
@@
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{1},\dots,a_{u},m_{u})\right)$ $=\sum_{t+t^{\prime}=nq_{u},t^{\prime}>0}\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),\left((n)\diamond({\underline{q}})\right)\circ_{j}(t,t^{\prime})}(a_{1},\dots,a_{u},m_{u}).$
@@
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{1},\dots,a_{u},m_{u})\right)=$ $\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma}x(y^{\times n}),(\lambda,\lambda^{\prime})\diamond({\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime}))}(a_{1},\dots,a_{u},a_{1},\dots,a_{u},m_{u}),$
@@
$(R_{\lambda,\lambda^{\prime}}^{l,l^{\prime}})_{u}=\left\{k-q_{u}+\alpha k+\gamma,\alpha\in[\lambda],\gamma\in[q_{u}]\right\}\cup\left\{\lambda k+(k-q_{u})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l]\right\}$
$(R_{\lambda,\lambda^{\prime}}^{l,l^{\prime}})_{u+1}=\left\{\lambda k+(k-l^{\prime})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l^{\prime}]\right\}.$
@@
$\sum_{\sigma^{\prime\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\Sigma_{(\lambda,\lambda^{\prime})\diamond({\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime}))}}\sigma\sum_{\Sigma_{(\lambda,\lambda^{\prime})\diamond({\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime}))}/\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma x(y^{\times n})$
$=\sum_{\sigma^{\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma^{\prime}x(y^{\times n}),$
@@
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{1},\dots,a_{u},m_{u})\right)$
$=\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma^{\prime}}\sigma^{\prime}x(y^{\times n}),R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}(a_{1},\dots,a_{u},m_{u}).$
@@
Assigning to each $A$-module $M$ the Beck module $A\ltimes M$ obtained by equipping the vector space $A\oplus M$ with the divided power $\mathcal{P}$-algebra structure given by (5.eq.1) provides a functor:
@@
There is an equivalence of categories between the category of abelian group objects in $\Gamma(\mathcal{P})$-algebras and the category of abelian $\Gamma(\mathcal{P})$-algebras in the sense of Definition 4.1, that is:
@@
Applying Theorem 5.2 in the case $A=0$, the data of an abelian group object in $\Gamma(\mathcal{P})$-algebras is equivalent to the data of a module over the $\Gamma(\mathcal{P})$-algebra 00. Such a 00-module $M$ is an abelian $\Gamma(\mathcal{P})$-algebra equipped with operations $\beta_{x,(r_{1},\dots,r_{s})}(0,\dots,0,-)\colon M\to M$. If $s>1$ and if at least one of the $r_{i}$ for $i<s$ is non-zero, then relation ($\beta$A3) implies that $\beta_{x,(r_{1},\dots,r_{s})}(0,\dots,0,m)=0$ for all $M$. If $s>1$ and $r_{i}=0$ for all $i<s$, relation ($\beta$A2) shows that $\beta_{x,(r_{1},\dots,r_{s})}(0,\dots,0,m)=\beta_{x,(r_{s})}(m)$. Finally, when $s=1$, relation ($\beta$AM7) shows that $\beta_{x,(r_{s})}(m)=\beta_{x}(m)$ is provided by the structure of abelian $\Gamma(\mathcal{P})$-algebra on $M$. Thus, a 00-module is an abelian $\Gamma(\mathcal{P})$-algebra equipped with additionnal operations which are all trivial. ∎
@@ 6.   Universal enveloping algebra
    $\beta_{\lambda x+y,{\underline{r}}}(a_{1},\dots,a_{s-1},-)-\left(\lambda\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)+\beta_{y,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\right)$ for all $x,y\in\mathcal{P}(n)^{{\underline{r}}}$, $\lambda\in\mathbb{F}$.
@@
    where ${\underline{q}}^{\prime}=({\underline{q}},0)$.
@@
    $\beta_{x,{\underline{r}}}(a_{0}+a_{1},\dots,a_{s-1},-)-\left(\sum_{l+l^{\prime}=r_{1}}\beta_{x,{\underline{r}}\circ_{1}(l,l^{\prime})}(a_{0},a_{1},\dots,a_{s-1},-)\right)$.
@@
    $\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)$ for all $x$ such that there exists $q_{1}+\dots+q_{u}=r_{s}$, $y\in\mathcal{P}(n)^{\Sigma_{{\underline{r}}\circ_{s}{\underline{q}}}}$, such that $x=\sum_{\sigma\in\Sigma_{n-r_{s}}\times\Sigma_{r_{s}}/\Sigma_{n-r_{s}}\times\Sigma_{{\underline{q}}}}\sigma y$, and at least two of the $q_{j}$’s are non-zero.
@@
Following Notation 4.6, we will use the notation:
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},-,a_{i+1},\dots,a_{s})\coloneqq\beta_{\sigma_{i}^{*},{\underline{r}}^{\sigma_{i}}}(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{s},-).$
@@
$\mu\colon\mathbb{U}_{\Gamma(\mathcal{P})}(A){}_{\mathbb{F}}\!\otimes_{\mathbb{F}}\mathbb{U}_{\Gamma(\mathcal{P})}(A)\to\mathbb{U}_{\Gamma(\mathcal{P})}(A)$
@@
The multiplication $\mu$ endows $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ with the structure of a unital ring.
@@
Let $V$ be the $1$-dimensional $\mathbb{F}$-vector space spanned by $v$. Consider the $\Gamma(\mathcal{P})$-algebra $B$ obtained as a quotient of the free $\Gamma(\mathcal{P})$-algebra $\Gamma(\mathcal{P},A\oplus V)$ by the $\Gamma(\mathcal{P})$-ideal generated by the two following family of elements:
@@
    where $z=x\left(1_{\mathcal{P}}^{\times n-r_{s}},y^{\times r_{s}}\right)\in\mathcal{P}(n+r_{s}(k-1))$, $R_{\lambda\lambda^{\prime}}^{ll^{\prime}}$ is defined in Definition 4.5, where $\tau$ ranges over $\Sigma_{{\underline{r}}\diamond_{s}{\underline{q}}}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{r_{s}}\wr\Sigma_{{\underline{q}}}$ in the sum, and where $\sigma^{\prime}$ ranges over $\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}$ in the sum.
@@
A concrete example is given by the enveloping algebra $V(A)=\mathbb{U}_{\Gamma(\mathrm{Com})}(A)$ studied in Section 9. In the special case $\mathbb{F}=\mathbb{F}_{p}$, the equality $\lambda^{r_{s}}=\lambda$ holds (since $r_{s}$ is a power of $p$), and thus $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ is an $\mathbb{F}$-algebra.
@@
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{j-1},-,a_{j+1},\dots,a_{s})\coloneqq\beta_{\sigma_{j}x,{\underline{r}}^{\sigma_{j}^{*}}}(a_{1},\dots,a_{j-1},a_{j+1},\dots,a_{s},-)$
in $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$.
@@
Denote by $\mathbb{U}_{\mathrm{Ab}}(\Gamma(\mathcal{P}))\subset\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ the vector subspace spanned by the elements $\beta_{x,(n)}(-)$. Note that $\mathbb{U}_{\mathrm{Ab}}(\Gamma(\mathcal{P}))$ is a unital subring and does not depend on $A$.
@@
$\beta_{x}(m)\coloneqq\beta_{x,(n)}(-)\cdot m.$
@@
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},m)\coloneqq\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\cdot m.$
@@
Let $M$ be an abelian $\Gamma(\mathcal{P})$-algebra. Then
$\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes_{\mathbb{U}_{\mathrm{Ab}(\Gamma(\mathcal{P}))}}M$
is the free $A$-module generated by $M$.
@@
Note that here, in the expression $\mathbb{U}_{\mathcal{P}}(A)$, we have omitted the forgetful functor $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\mathcal{P}\text{-}\mathrm{Alg}$ which is induced by the norm map as well.